# Spatialyze: A Geospatial Video Analytics System with Spatial-Aware Optimizations

### v4

@@ 1. Introduction
Leveraging this insight, we present Spatialyze, a system for geospatial video analytics. To make it easy for users to specify their geospatial video analytics workflows, Spatialyze comes with a conceptual data model where users create and ingest videos into a “world,” and users interact with the world by specifying objects (e.g., cars) and scenarios (e.g., cars at an intersection) of interest via Spatialyze’s S-Flow, a domain-specific language (DSL) embedded in Python. Spatialyze then efficiently executes the workflow by leveraging various spatial-aware optimization techniques that make use of existing geospatial metadata and assumptions based on the inherited physical behavior of objects in the videos. For instance, Spatialyze’s Road Visibility Pruner uses the road’s visibility as a proxy for objects’ visibility to prune out video frames. The Object Type Pruner then prunes out objects that are not of interest to the users’ workflow. Spatialyze’s 0pt speeds up object 3D location estimation by replacing a computationally expensive ML-based approach with a geometry-based approach. Finally, the Exit Frame Sampler prunes out unnecessary video frames based on the inherited physical behavior of vehicles and traffic rules. Spatialyze’s optimizations are all driven by various geospatial metadata and real-world physical behavior, with the goal of reducing the number of video frames to be processed and ML operations to be invoked, thus speeding up the video processing runtime.
@@ 3. Usage Scenario
Spatialyze streamlines this process. Instead of manually writing code to run ML models and extract geospatial information, journalists can use Spatialyze to construct such workflows easily. With Spatialyze, they can construct an analytic workflow with S-Flow’s 3-step paradigm as shown in § 3 to: Build a world, by integrating the video data with the geospatial metadata (e.g., road network and camera configurations); Filter the video for parts of interest, by describing relationships between objects (humans, cars, trucks) and their surrounding geospatial environment (lane, intersection); Observe the filtered video parts, in this case, by saving all the filtered video parts into video files for further examination.
Build a world. In line 3, the user creates a world which represents a geospatial virtual environment to be discussed in § 4. Line 3 then initializes the world by loading roadNetwork data, in this case, a set of the Boston Seaport’s road segments. Lines 3-3 load videos with their camera configurations (locations, rotations, and intrinsics (Zhang, 2014)) and associate them to create geospatial videos added into the world. Finally, in line 3, they have constructed a data-rich world with video cameras, and road segments, all related spatially and temporally.
Filter for the video parts of interest. From the data journalist’s perspective, they can interact with objects and their trackings from each video added to the world through object() with the world initialized. Here, our journalist wants to find video snippets where other vehicles crash into the driver’s vehicle at a road intersection. They do so by adding a filter to the world to only include objects that match the provided predicates: that they are cars or trucks (line 3), are within 50 meters of the camera (line 3), are at an intersection (line 3), and are moving towards the camera (in line 3). After filter is executed at line 3, the world will contain only cars or trucks on intersections that are moving towards the camera, as desired.
Observe the filtered video parts. After applying the filter, the world has fewer objects than when it was initialized. When the journalist calls saveVideos() in line 3, Spatialyze saves only video snippets with objects in the current world, significantly shortening the total video length. This allows the journalist to focus only on relevant segments with highlighted objects of interest. As shown, S-Flow provides a declarative interface where users do not specify “how” to find the video snippets of interest; instead, they only need to describe “what” their video snippets of interest look like, and Spatialyze will optimize accordingly.
@@
We next discuss Spatialyze in detail (in § 5) and how it optimizes workflow execution (in § 6) by leveraging videos’ geospatial metadata and the physical behavior of real-world objects. {listing}[!ht]{minted}[ linenos, mathescape, escapeinside=!!, fontsize=]python world = World() # OR World(detect=CstmDetector, track=CstmTracker) world.addGeogConstructs(RoadNetwork(’road-network/boston-seaport/’)) !! world.addVideo(GeospatialVideo(Video(’v0.mp4’), Camera(’c0.json’))) !! world.addVideo(GeospatialVideo(Video(’v1.mp4’), Camera(’c1.json’))) !! obj, cam, intersection = ( world.object(), world.camera(), !! world.geogConstruct(type=’intersection’) ) world.filter( ((obj.type == ’car’) — (obj.type == ’truck’)) !! !! & (distance(obj, cam) ¡ 50) contains(intersection, obj) !! !! headingDiff(obj, cam, between=[135, 225]) ) !! world.saveVideos(’output_videos/’, addBoundingBoxes=True) !! Geospatial Video Analytic Workflow with Spatialyze
@@ 4.1.1. World
As discussed in § 3 and in line 3 of § 3, users construct a World as a first step of S-Flow. Inspired by VisualWorldDB, a World represents a geospatial virtual environment which encompasses Geographic Constructs and Movable Objects, all coexisting with spatial relationships. These relationships, combined with the S-Flow, enable users to describe the videos of interest effectively.
@@ 4.2.1. Camera
Lines 3-3 of § 3 defines a Movable Object with $type=camera$ as discussed in in § 4.1.3. Users initialize a Camera from a list, where the $i$-th element of the list is a dictionary with 4 fields: translation, rotation, intrinsic, and timestamp. They correspond to $l_{i}$, $r_{i}$, $it_{i}$, and $t_{i}$, as defined in § 4.1.3.
@@ 4.2.2. GeospatialVideo
A GeospatialVideo is a video enhanced with geospatial metadata. A Video (lines 3-3 of § 3) by itself contains a sequence of images where each represents a frame of the video. Such video exists spatially in the World after its Video object is bound to a Camera to obtain its spatial properties, where each frame $i$ is an image taken from the camera at location $l_{i}$, rotation $r_{i}$, intrinsic $it_{i}$, and at time $t_{i}$. An object in each frame obtains its geospatial properties after its track is inferred from GeospatialVideo. These objects then become Movable Objects as defined in § 4.1.3.
@@ 4.2.4. World
Build. A World is initially empty. Users add RoadNetwork using addGeogContructs(), followed by adding GeospatialVideo via addVideo(). At this point of the workflow, the World contains Geographic Constructs, and Cameras. Spatialyze uses YOLOv5 (Jocher et al., 2022) and StrongSORT (Du et al., 2023) for object detection and object tracking by default. Spatialyze also supports using UDFs for object detection or tracking, shown as the comment in line 3 of § 3.
@@ 5. Workflow Execution in Spatialyze
[…]
Spatialyze System Image: geospatial video analytic workflow execution. The user writes an S-Flow workflow in Python along with the input videos, cameras, and road network. Spatialyze’s Data Integrator integrates the videos and cameras into geospatial videos and ingests the road network into the geospatial metadata store. Its Video Processor processes the geospatial videos to get the Movable Objects. Its Movable Objects Query Engine filters the relevant Movable Objects based on the user’s filter. Finally, its Output Composer composes either the videos with the relevant Movable Objects or Object Trajectories for the user.
Spatialyze optimizes a video processing plan by inserting optimization operators into the plan. The Road Visibility Pruner is before the Video Decoder. The Object Type Pruner is right after the Object Detector. The Geometry-Based 3D Location Estimator replaces the ML-based 3D Location Estimator. The Exit Frame Sampler is right before the Object Tracker.
[…]
From users’ view, Spatialyze filters objects in the World after they call filter(). Then, they can observe the filtered objects afterward, using one of the observer functions described in § 4.2.4. Internally, Spatialyze only performs video processing to detect objects and compute tracks for them as needed given the workflow. For example, in line 3 of § 3, the user filters on objects’ moving directions. Executing this requires computing the trajectory of objects in the videos to compute their directions. Given an S-Flow workflow, the goal of Spatialyze is to understand the users’ workflow, recognize and execute only the video processing operators needed for users’ filter predicates, and return Movable Objects that satisfy the filters.
@@ 5.2. System Design & Workflow Execution
[…]
@@ 5.2.2. Video Processor
[…]
@@
{listing}
[h!]{minted}[linenos,mathescape,escapeinside=!!,fontsize=]python frames = VideoDecoder(’videofile.mp4’) !! object2Ds = ObjectDetector(frames) !! object2DAnd3Ds = Loc3DEstm(frames, object2Ds, cameras) !! movableObjects = ObjectTracker(object2DAnd3Ds, frames) !! Baseline Video Processing Plan
Executing the plan computes 3D tracks of objects in a video. However, streaming each video through all operators is computationally expensive and not always necessary. Based on the users’ filter predicates from when they Filter the World, Spatialyze only includes the necessary operators in the execution plan. For instance, a predicate based on object types like obj.type=="car" requires objects’ types and hence the execution plan will include lines 5.2.2 and 5.2.2. Likewise, a predicate with distance or contain functions requires objects’ 3D location and therefore include lines 5.2.2, 5.2.2, and 5.2.2 in the plan. Finally, a predicate with a headingDiff function requires objects’ moving directions and include all lines in the plan.
@@ 5.2.4. Output Composer
Fig. 1 shows the execution stages for the workflow in § 3. It contains all the stages mentioned above. A user first builds the world by adding data. Spatialyze’s Data Integrator creates an indexed table consisting of RoadNetwork and joins the Video and its corresponding Camera. Then, Spatialyze’s Video Processor creates an optimized video processing plan with 4 processing operators and 4 optimization operators (to be discussed in § 6). Parts of the users’ predicates are executed in Road Visibility Pruner and Object Type Pruner. The Movable Objects Query Engine streams the Movable Objects from the Video Processor into the geospatial metadata store and filters them with the users’ predicates. Finally, as the workflow asks for videos, Spatialyze formats and saves the frames that have the filtered objects into video snippet files.
@@ 6. Video Processing Optimization
SOTA video inference uses ML models to extract the desired information, such as objects’ positions and trajectories within each video. Running such models is computationally expensive. Spatialyze’s video processor operates with 4 optimization techniques to reduce such runtime bottlenecks. To be discussed below, the idea behind our optimization is to leverage the geospatial properties of the videos, and the Movable Objects’ inherited physical behaviors. We implement each optimization as a streaming operator in the Video Processing stage in § 5.2.2. In addition to the existing 4 processing operators, Spatialyze may place optimization operators in the middle of the video processing plan to reduce the amount of input into later processing operators based on the filter predicates. In § 6.1, Spatialyze places the Road Visibility Pruner after the Video Decoder, when users specify the Geographic Constructs the objects of interest must be in. Doing so reduces the number of input frames for the rest of the operators in the plan. In § 6.2, If users filter on object types, then Spatialyze places the Object Type Pruner right after the Object Detector. It reduces the number of objects that Spatialyze needs to calculate for their 3D locations and tracking. In § 6.3, Spatialyze replaces the ML-based 3D Location Estimation operator with the much faster 0pt, if Spatialyze can assume that the objects of interest are on the ground, such as cars or people. In § 6.4, Spatialyze places the Exit Frame Sampler in between the 3D Location Estimator and the Object Tracker, if the users’ predicates focus on cars. It utilizes the 3D locations of the objects and the user-provided metadata to reduce the number of frames that the Object Tracker processes.
Fig. 1 shows where the video processor executes all the optimization steps. Line 3 in § 3 corresponds to Road Visibility Pruner. Line 3 filters by object type, so Spatialyze adds Object Type Pruner. The objects of interest are cars or trucks, so Spatialyze adds Exit Frame Sampler and 0pt. In the following sections, we discuss each optimization in detail.
@@ 6.1.1. High-Level Concepts
The Road Visibility Pruner uses Geographic Construct’s visibility as a proxy for a Movable Object’s visibility. Specifically, a predicate includes contains(road,obj) and distance(cam,obj)<d means that obj is not visible in the frames where road is not visible within d meters; therefore, Spatialyze prunes out those frames. In § 3, the Road Visibility Pruner partially executes lines 3 where the user searches for “cars within 50 meters at an intersection.” Spatialyze removes all frames that do not contain any visible intersection within 50 meters as these frames will not contain any car of users’ interest.
@@ 6.1.2. Algorithm
[…]
@@
(3) $\begin{split}\left[\begin{smallmatrix}x_{c}\\ y_{c}\\ z_{c}\\ 1\end{smallmatrix}\right]&=\left[\begin{smallmatrix}\begin{array}[]{c|c}\text{\small\emph{it}}_{\text{\scriptsize\em i}}^{\text{\scriptsize-1}}&\text{\small{0}}\\ \hline\cr\text{\small{0}}&\text{\small{1}}\end{array}\end{smallmatrix}\right]\times\left[\begin{smallmatrix}x_{p}z_{c}\\ y_{p}z_{c}\\ z_{c}\\ 1\end{smallmatrix}\right]\\[-8.0pt] &=C\times\left[\begin{smallmatrix}x_{p}z_{c}&y_{p}z_{c}&z_{c}&1\end{smallmatrix}\right]^{\top},\text{where }C=\left[\begin{smallmatrix}\begin{array}[]{c|c}\text{\small\emph{it}}_{\text{\scriptsize\em i}}^{\text{\scriptsize-1}}&\text{\small{0}}\\ \hline\cr\text{\small{0}}&\text{\small{1}}\end{array}\end{smallmatrix}\right]\end{split}$
@@
[…]

### v5

@@ 1. Introduction
Leveraging this insight, we present Spatialyze, a system for geospatial video analytics. To make it easy for users to specify their geospatial video analytics workflows, Spatialyze comes with a conceptual data model where users create and ingest videos into a “world,” and users interact with the world by specifying objects (e.g., cars) and scenarios (e.g., cars at an intersection) of interest via Spatialyze’s S-Flow, a domain-specific language (DSL) embedded in Python. Spatialyze then efficiently executes the workflow by leveraging various spatial-aware optimization techniques that make use of existing geospatial metadata and assumptions based on the inherited physical behavior of objects in the videos. For instance, Spatialyze’s Road Visibility Pruner uses the road’s visibility as a proxy for objects’ visibility to prune out video frames. The Object Type Pruner then prunes out objects that are not of interest to the users’ workflow. Spatialyze’s Geometry-Based 3D Location Estimator speeds up object 3D location estimation by replacing a computationally expensive ML-based approach with a geometry-based approach. Finally, the Exit Frame Sampler prunes out unnecessary video frames based on the inherited physical behavior of vehicles and traffic rules. Spatialyze’s optimizations are all driven by various geospatial metadata and real-world physical behavior, with the goal of reducing the number of video frames to be processed and ML operations to be invoked, thus speeding up the video processing runtime.
@@ 3. Usage Scenario
Spatialyze streamlines this process. Instead of manually writing code to run ML models and extract geospatial information, journalists can use Spatialyze to construct such workflows easily. With Spatialyze, they can construct an analytic workflow with S-Flow’s 3-step paradigm as shown in 1 to: Build a world, by integrating the video data with the geospatial metadata (e.g., road network and camera configurations); Filter the video for parts of interest, by describing relationships between objects (humans, cars, trucks) and their surrounding geospatial environment (lane, intersection); Observe the filtered video parts, in this case, by saving all the filtered video parts into video files for further examination.
Build a world. In line 1, the user creates a world which represents a geospatial virtual environment to be discussed in § 4. Line 2 then initializes the world by loading roadNetwork data, in this case, a set of the Boston Seaport’s road segments. Lines 3-4 load videos with their camera configurations (locations, rotations, and intrinsics (Zhang, 2014)) and associate them to create geospatial videos added into the world. Finally, in line 4, they have constructed a data-rich world with video cameras, and road segments, all related spatially and temporally.
Filter for the video parts of interest. From the data journalist’s perspective, they can interact with objects and their trackings from each video added to the world through object() with the world initialized. Here, our journalist wants to find video snippets where other vehicles crash into the driver’s vehicle at a road intersection. They do so by adding a filter to the world to only include objects that match the provided predicates: that they are cars or trucks (line 7), are within 50 meters of the camera (line 8), are at an intersection (line 8), and are moving towards the camera (in line 9). After filter is executed at line 9, the world will contain only cars or trucks on intersections that are moving towards the camera, as desired.
Observe the filtered video parts. After applying the filter, the world has fewer objects than when it was initialized. When the journalist calls saveVideos() in line 10, Spatialyze saves only video snippets with objects in the current world, significantly shortening the total video length. This allows the journalist to focus only on relevant segments with highlighted objects of interest. As shown, S-Flow provides a declarative interface where users do not specify “how” to find the video snippets of interest; instead, they only need to describe “what” their video snippets of interest look like, and Spatialyze will optimize accordingly.
@@
We next discuss Spatialyze in detail (in § 5) and how it optimizes workflow execution (in § 6) by leveraging videos’ geospatial metadata and the physical behavior of real-world objects.
Listing 1: Geospatial Video Analytic Workflow with Spatialyze
1world = World() # OR World(detect=CstmDetector, track=CstmTracker)
2world.addGeogConstructs(RoadNetwork(’road-network/boston-seaport/’))
3world.addVideo(GeospatialVideo(Video(’v0.mp4’), Camera(’c0.json’)))
4world.addVideo(GeospatialVideo(Video(’v1.mp4’), Camera(’c1.json’)))
5obj, cam, intersection = ( world.object(), world.camera(),
6 world.geogConstruct(type=’intersection’) )
7world.filter( ((obj.type == ’car’) | (obj.type == ’truck’))
8 & (distance(obj, cam) < 50) & contains(intersection, obj)
9 & headingDiff(obj, cam, between=[135, 225]) )
10world.saveVideos(’output_videos/’, addBoundingBoxes=True)
@@ 4.1.1. World
As discussed in § 3 and in line 1 of 1, users construct a World as a first step of S-Flow. Inspired by VisualWorldDB, a World represents a geospatial virtual environment which encompasses Geographic Constructs and Movable Objects, all coexisting with spatial relationships. These relationships, combined with the S-Flow, enable users to describe the videos of interest effectively.
@@ 4.2.1. Camera
Lines 3-4 of 1 defines a Movable Object with $type=camera$ as discussed in in § 4.1.3. Users initialize a Camera from a list, where the $i$-th element of the list is a dictionary with 4 fields: translation, rotation, intrinsic, and timestamp. They correspond to $l_{i}$, $r_{i}$, $it_{i}$, and $t_{i}$, as defined in § 4.1.3.
@@ 4.2.2. GeospatialVideo
A GeospatialVideo is a video enhanced with geospatial metadata. A Video (lines 3-4 of 1) by itself contains a sequence of images where each represents a frame of the video. Such video exists spatially in the World after its Video object is bound to a Camera to obtain its spatial properties, where each frame $i$ is an image taken from the camera at location $l_{i}$, rotation $r_{i}$, intrinsic $it_{i}$, and at time $t_{i}$. An object in each frame obtains its geospatial properties after its track is inferred from GeospatialVideo. These objects then become Movable Objects as defined in § 4.1.3.
@@ 4.2.4. World
Build. A World is initially empty. Users add RoadNetwork using addGeogContructs(), followed by adding GeospatialVideo via addVideo(). At this point of the workflow, the World contains Geographic Constructs, and Cameras. Spatialyze uses YOLOv5 (Jocher et al., 2022) and StrongSORT (Du et al., 2023) for object detection and object tracking by default. Spatialyze also supports using UDFs for object detection or tracking, shown as the comment in line 1 of 1.
@@ 5. Workflow Execution in Spatialyze
Figure 1. (a) Geospatial Video Analytics Workflow Execution. (b) Processing Plan for the workflow shown in 1.
From users’ view, Spatialyze filters objects in the World after they call filter(). Then, they can observe the filtered objects afterward, using one of the observer functions described in § 4.2.4. Internally, Spatialyze only performs video processing to detect objects and compute tracks for them as needed given the workflow. For example, in line 9 of 1, the user filters on objects’ moving directions. Executing this requires computing the trajectory of objects in the videos to compute their directions. Given an S-Flow workflow, the goal of Spatialyze is to understand the users’ workflow, recognize and execute only the video processing operators needed for users’ filter predicates, and return Movable Objects that satisfy the filters.
@@ 5.2. System Design & Workflow Execution
To minimize execution time, Spatialyze defers ML inferences, which are expensive, until users observe the World. This way, Spatialyze can optimize the entire workflow execution. Once users observe a World, Spatialyze starts executing users’ workflow in 4 stages sequentially: (1) Data Integrator, (2) Video Processor, (3) Movable Objects Query Engine, and (4) Output Composer; as shown in Fig. 1.
@@ 5.2.2. Video Processor
We designed the video processor to stream video frames through all the above algorithms. Each video frame is pipelined through decoding, object detection, 3D location estimation, and object tracking; hence our video processor only needs to keep $O(1)$ video frames and can handle videos of arbitrary size, addressing the second challenge. While stream processing of videos has been used previously (Broström, 2022), applying it to a streaming video processor is not trivial. For example, the original YOLO-StrongSORT object tracker (Broström, 2022) streams video frames. However, naively extending the tracker with our optimization techniques requires hard-coding them into the tracker’s implementation. This does not scale to future trackers. To solve this problem (and address the third challenge), we implemented our video processor as a plan of streaming operators shown in 2. Our baseline operators111 The streaming operators internally process videos and geospatial metadata to extract (2D or 3D) object detections and/or tracks. In contrast, users use predicate operators (e.g., contains, distance) to define filter predicates. include: (1) Video Decoder; (2) Object Detector; (3) 3D Location Estimator; and (4) Object Tracker. Specifically, (1) decodes each video to get a sequence of RGB images representing video frames. (2) detects objects in each frame from (1), returning each object and its object type and 2D bounding box. AWorld is 3 dimensional where all objects have 3D locations; therefore, (3) estimates the 3D location of each object from (2) based on its Camera. Lastly, (4) tracks objects by associating detections from (3) between frames. We implemented each operator as an iterator function, where each operator takes in streams of per-frame inputs and returns a stream of per-frame outputs. As a result, our video processor can process videos of arbitrary size, including those larger than the available memory. In addition, we implemented a caching system for the operators. In 2, frames is an input to both ObjectDetector and ObjectTracker. Instead of computing the stream of frames twice, Spatialyze iterates through frames once and caches the results. For instance, a frame $i$ is decoded when the ObjectDetector processes it. Spatialyze caches the frame $i$ so that the ObjectTracker can process it without having to decode the frame again. Spatialyze keeps track of the number of the operators that will be using a cached result and evicts a frame as soon as all of the operators finished processing it; therefore, Spatialyze only caches a few frames at a given time.
@@
Listing 2: Baseline Video Processing Plan
1frames = VideoDecoder(’videofile.mp4’)
2object2Ds = ObjectDetector(frames)
3object2DAnd3Ds = Loc3DEstm(frames, object2Ds, cameras)
4movableObjects = ObjectTracker(object2DAnd3Ds, frames)
Executing the plan computes 3D tracks of objects in a video. However, streaming each video through all operators is computationally expensive and not always necessary. Based on the users’ filter predicates from when they Filter the World, Spatialyze only includes the necessary operators in the execution plan. For instance, a predicate based on object types like obj.type=="car" requires objects’ types and hence the execution plan will include lines 1 and 2. Likewise, a predicate with distance or contain functions requires objects’ 3D location and therefore include lines 1, 2, and 3 in the plan. Finally, a predicate with a headingDiff function requires objects’ moving directions and include all lines in the plan.
@@ 5.2.4. Output Composer
Fig. 1 shows the execution stages for the workflow in 1. It contains all the stages mentioned above. A user first builds the world by adding data. Spatialyze’s Data Integrator creates an indexed table consisting of RoadNetwork and joins the Video and its corresponding Camera. Then, Spatialyze’s Video Processor creates an optimized video processing plan with 4 processing operators and 4 optimization operators (to be discussed in § 6). Parts of the users’ predicates are executed in Road Visibility Pruner and Object Type Pruner. The Movable Objects Query Engine streams the Movable Objects from the Video Processor into the geospatial metadata store and filters them with the users’ predicates. Finally, as the workflow asks for videos, Spatialyze formats and saves the frames that have the filtered objects into video snippet files.
@@ 6. Video Processing Optimization
SOTA video inference uses ML models to extract the desired information, such as objects’ positions and trajectories within each video. Running such models is computationally expensive. Spatialyze’s video processor operates with 4 optimization techniques to reduce such runtime bottlenecks. To be discussed below, the idea behind our optimization is to leverage the geospatial properties of the videos, and the Movable Objects’ inherited physical behaviors. We implement each optimization as a streaming operator in the Video Processing stage in § 5.2.2. In addition to the existing 4 processing operators, Spatialyze may place optimization operators in the middle of the video processing plan to reduce the amount of input into later processing operators based on the filter predicates. In § 6.1, Spatialyze places the Road Visibility Pruner after the Video Decoder, when users specify the Geographic Constructs the objects of interest must be in. Doing so reduces the number of input frames for the rest of the operators in the plan. In § 6.2, If users filter on object types, then Spatialyze places the Object Type Pruner right after the Object Detector. It reduces the number of objects that Spatialyze needs to calculate for their 3D locations and tracking. In § 6.3, Spatialyze replaces the ML-based 3D Location Estimation operator with the much faster Geometry-Based 3D Location Estimator, if Spatialyze can assume that the objects of interest are on the ground, such as cars or people. In § 6.4, Spatialyze places the Exit Frame Sampler in between the 3D Location Estimator and the Object Tracker, if the users’ predicates focus on cars. It utilizes the 3D locations of the objects and the user-provided metadata to reduce the number of frames that the Object Tracker processes.
Fig. 1 shows where the video processor executes all the optimization steps. Line 8 in 1 corresponds to Road Visibility Pruner. Line 7 filters by object type, so Spatialyze adds Object Type Pruner. The objects of interest are cars or trucks, so Spatialyze adds Exit Frame Sampler and Geometry-Based 3D Location Estimator. In the following sections, we discuss each optimization in detail.
@@ 6.1.1. High-Level Concepts
The Road Visibility Pruner uses Geographic Construct’s visibility as a proxy for a Movable Object’s visibility. Specifically, a predicate includes contains(road,obj) and distance(cam,obj)<d means that obj is not visible in the frames where road is not visible within d meters; therefore, Spatialyze prunes out those frames. In 1, the Road Visibility Pruner partially executes lines 8 where the user searches for “cars within 50 meters at an intersection.” Spatialyze removes all frames that do not contain any visible intersection within 50 meters as these frames will not contain any car of users’ interest.
@@ 6.1.2. Algorithm
First, the Road Visibility Pruner calculates the 3D viewable space of the Camera at frame $i$. The viewable space is a pyramid (Fig. 2 side view). (A) is the location of the camera. (B), (C), (D), and (E) represent the top-left, top-right, bottom-right, and bottom-left of the frame at $d$ meters in front of the camera. We compute (B), (C), (D), and (E) by converting the mentioned 4 corner points of the frame from pixel (2D) to world (3D) coordinates. This is done in 3 steps by converting a point from: 1) Pixel to Camera Coordinate System; 2) Camera to World Coordinate System; and 3) Pixel to World Coordinate System.
@@
(3) $\begin{split}\left[\begin{smallmatrix}x_{c}\\ y_{c}\\ z_{c}\\ 1\end{smallmatrix}\right]&=\left[\begin{smallmatrix}\begin{array}[]{c|c}\text{\small\emph{\text{$it_{i}^{-1}$}}}&\text{\small{0}}\\ \hline\cr\text{\small{0}}&\text{\small{1}}\end{array}\end{smallmatrix}\right]\times\left[\begin{smallmatrix}x_{p}z_{c}\\ y_{p}z_{c}\\ z_{c}\\ 1\end{smallmatrix}\right]\\[-8.0pt] &=C\times\left[\begin{smallmatrix}x_{p}z_{c}&y_{p}z_{c}&z_{c}&1\end{smallmatrix}\right]^{\top},\text{where }C=\left[\begin{smallmatrix}\begin{array}[]{c|c}\text{\small\emph{\text{$it_{i}^{-1}$}}}&\text{\small{0}}\\ \hline\cr\text{\small{0}}&\text{\small{1}}\end{array}\end{smallmatrix}\right]\end{split}$
@@
After the corners of each frame are translated to world coordinates, the Road Visibility Pruner uses the viewable space to determine the Geographic Constructs that are visible by Camera at frame $i$. Each Geographic Construct’s area is defined using a 2D polygon. Specifically, all polygons representing Geographic Constructs are 2D in the $z=0$ plane. Therefore, the Road Visibility Pruner projects the computed 3D viewable space onto the $z=0$ plane. We define the top-down 2D viewable area as the convex hull (Sklansky, 1982) of the projected vertices (A), (B), (C), (D), and (E) (yellow highlight in Fig. 2’s top view). We then spatially join the 2D viewable area with Geographic Constructs in the geospatial metadata store, using their geospatial indexing created in § 5.2.1. Any Geographic Construct that overlaps with the 2D viewable area is considered visible to the Camera at frame $i$. Once computed, we get visibleGeogs, a set of Geographic Construct types that are visible to the Camera at frame $i$.